# Sara Najem

### v1

@@ Abstract
Determining whether two graphs are structurally identical is a fundamental problem with applications spanning mathematics, computer science, chemistry, and network science. Despite decades of study, graph isomorphism remains a challenging algorithmic task, particularly for highly symmetric structures. Here we introduce a new algorithmic approach based on ideas from spectral graph theory and geometry that constructs candidate correspondences between vertices using their curvatures. Any correspondence produced by the algorithm is explicitly verified, ensuring that non-isomorphic graphs are never incorrectly identified as isomorphic. Although the method does not yet guarantee success on all isomorphic inputs, we find that it correctly resolves every instance tested in deterministic polynomial time, including a broad collection of graphs known to be difficult for classical spectral techniques. These results demonstrate that enriched spectral methods can be far more powerful than previously understood, and suggest a promising direction for the practical resolution of the complexity of the graph isomorphism problem.
@@ 1 Introduction
In this work, we introduce a new algorithmic approach to graph isomorphism based on spectral geometry. Rather than exploiting combinatorial restrictions, our method draws on ideas from non-commutative geometry and spectral graph theory, using curvature derived from Laplacian spectra as a tool for symmetry breaking. The conceptual foundation of our approach is inspired by recent work on geometric features of networks based on Connes’ spectral triplet formalism [14], where distance, dimension, and curvature are recovered from spectral data via heat kernel methods. While that framework was developed for higher-order networks and simplicial complexes, we adapt its core ideas to the setting of simple graphs and graph isomorphism.
Our algorithm assigns to each vertex a curvature value computed from the short-time behavior of the heat kernel associated with the graph Laplacian. These curvature values are then used to construct a candidate correspondence between vertices of two graphs. Any correspondence produced by the algorithm is explicitly verified, ensuring that the method never falsely declares two non-isomorphic graphs to be isomorphic. Consequently, the algorithm is one-sidedly correct: it always identifies no-instances correctly, and any potential failure can only occur on isomorphic inputs.
We evaluate the algorithm on an extensive collection of graphs generated using standard generators, encompassing a wide range of sizes, densities, and structural properties. Remarkably, the method successfully resolves all instances tested, including graphs that are highly symmetric and traditionally challenging for purely spectral approaches. While this empirical success does not constitute a proof of polynomial-time solvability of graph isomorphism, it provides strong evidence that curvature-based spectral information captures structural features that are invisible to classical eigenvalue methods.
Beyond its practical performance, this work establishes a previously unexplored connection between graph isomorphism and discrete spectral geometry. It demonstrates that geometric quantities derived from spectral data, specifically curvature, can serve as effective symmetry-breaking invariants. This suggests a new direction for GI algorithms, complementary to group-theoretic and combinatorial techniques, and raises the possibility that spectral geometry may play a substantive role in future progress toward polynomial-time graph isomorphism algorithms.
@@ 2.1 Spectral graph theory, Laplacians, and heat kernels
The graph Laplacian is a central object in spectral graph theory and provides a bridge between discrete graph structure and continuous analytic methods. Given an undirected graph $G=(V,E)$ with adjacency matrix $A$ and degree matrix $D$ (where $D_{ii}=\deg(i)$), the combinatorial Laplacian is defined as L=DAL=D-A. The matrix $L$ is symmetric and positive semidefinite, and its eigenvalues are real and nonnegative [13]. A fundamental property of $L$ is that it always admits the constant vector as an eigenvector with eigenvalue 0. More generally, the multiplicity of the 0 eigenvalue equals the number of connected components of $G$, with each component contributing one independent vector to the kernel of $L$ [6]. Throughout this work we focus on the combinatorial Laplacian; weighted graphs and normalized variants can be treated analogously but are not required for our analysis.
@@
These connections motivate the interpretation of $L$ as a discrete analogue of the Laplace–Beltrami operator on a Riemannian manifold. In this analogy, the heat kernel $H(t)=e^{-tL}$ plays a role analogous to its continuous counterpart. The heat kernel describes the evolution of diffusion on the graph and provides a multi-scale summary of its structure: small values of $t$ probe local neighborhoods, while larger $t$ reflect global connectivity. Heat-kernel-based constructions have found applications in clustering, graph comparison, and shape analysis, and they form the foundation for several modern spectral descriptors [7].
@@ 2.2 Spectral invariants and graph isomorphism
Despite this limitation, spectral methods play an important role in practical GI algorithms. Eigenvalues and eigenvectors are commonly used as preprocessing tools to partition vertices into equivalence classes, reduce symmetry, and guide subsequent combinatorial searches. More refined spectral constructions attempt to extract richer invariants from Laplacian eigenfunctions rather than relying solely on eigenvalues. For example, heat-kernel-based signatures embed vertices into continuous feature spaces that capture structural information across multiple scales [15]. Such methods transform isomorphism testing into a matching problem between spectral embeddings, often yielding efficient algorithms in practice even though worst-case guarantees remain elusive.
@@ 2.4 From spectra to geometry
Recent advances in spectral geometry and non-commutative geometry provide precisely such tools. By extracting geometric quantities, such as distances, dimensions, and curvature, from Laplacian spectra and heat kernels, one can associate a geometric structure to a graph that is invariant under relabeling but sensitive to subtle asymmetries. In this work, we build on these ideas to develop curvature-based spectral features for simple graphs and demonstrate their effectiveness for graph isomorphism testing.
@@ 3 The base algorithm
We will use Equations 2 and 5 to retrieve every node’s curvature in each graph where $u_{1}(i,i)$ will be computed from the polynomial fit of $K(i,i,t)$ versus $t$ up to $k=\text{degree(i)}$, which is the degree of node $i$. Recall that the expansion is only valid in the vicinity of $t^{+}\rightarrow 0$. Therefore, we first evaluate Equation 2 over the interval $t\in[t_{min}=1/\lambda_{max},t^{\prime}]$, where $\lambda_{max}$ is the largest eigenvalue of $L$. It defines the characteristic time scale for the heat to start diffusing over the network, while $t^{\prime}\leq 1/\lambda_{min}$, which is the time needed for the heat to diffuse over all the graphs in the network. Then a segmented linear fit of the $\log K(i,i,t)$ is applied on the data to detect a change in behavior of $K(i,i,t)$. The first point of change denotes $t_{max}$, the upper limit on the time interval over which the kernel should be followed for node $i$. Subsequently, we sample from $K(i,i,t)$, on the identified interval $t\in[t_{min}=1/\lambda_{max},t_{max}$, a number of points equal to $\text{degree}_{i}$ and fit them against the powers of $t$ given in Equation 5 to recover the curvatures.
@@
4 The generalized algorithm
The core algorithm described earlier relies on curvature values derived from spectral geometry to induce a vertex correspondence between two graphs. While this approach performs well in a wide range of instances, two types of degeneracies can arise in practice: graphs whose spectra are highly degenerate or insufficiently smooth, and graphs with large automorphism groups that preserve curvature values across many vertices. In this section, we describe a generalized version of the algorithm that addresses both issues while preserving equivalence.
@@ 4.1 Stretching geometry using edge subdivisions
4.2 Handling large automorphism groups via vertex identification
A second source of difficulty arises when graphs possess large automorphism groups. In such cases, many vertices may share identical curvature values even after geometric refinement, reflecting genuine symmetries of the graph. While this does not lead to false positives, it can prevent the algorithm from constructing a unique vertex correspondence.
To address this, we introduce an iterative vertex identification procedure that systematically breaks residual symmetries. The procedure operates as follows. At each iteration, we identify the smallest nontrivial equivalence class of vertices in the first graph whose members share the same curvature value. We perform the same identification in the second graph. If the sizes of these classes differ, the graphs cannot be isomorphic and the algorithm terminates. Otherwise, we select one representative vertex from each class.
To distinguish the chosen vertices without altering isomorphism status, we attach to each selected vertex a fixed number of disjoint triangles, denoted by an iteration-dependent parameter. Concretely, at iteration $t$, we attach $t$ new triangles to the selected vertex by adding $3t$ new vertices and $3t$ edges forming $t$ isolated 3-cycles incident only to that vertex. This operation creates a unique local geometric signature that is reflected immediately in curvature values, while remaining invariant under graph isomorphism when applied symmetrically.
The modified graphs are then reprocessed by the curvature-based algorithm, producing refined curvature values that typically split the previously degenerate equivalence classes. If necessary, the process is repeated, incrementing the number of attached triangles at each iteration. Because each step introduces strictly local modifications and preserves isomorphism, the procedure is guaranteed to maintain one-sided correctness: non-isomorphic graphs are never identified as isomorphic.
@@ 4.2 Handling large automorphism groups via vertex identification
Together, edge subdivision and vertex identification extend the applicability of the curvature-based approach to degenerate spectral regimes and graphs with large automorphism groups. These mechanisms allow the algorithm to operate robustly across a broader class of inputs while preserving correctness and maintaining a clear geometric interpretation.
@@ 5 Computational complexity
We analyze the running time of (i) the base curvature–matching algorithm and (ii) the generalized algorithm. Throughout, let $n=|V|$ and $m=|E|$ for the input graphs. We assume standard RAM-model arithmetic on $O(\log n)$-bit numbers and treat polynomial regression on $O(d)$ points as $O(d^{3})$ in the worst case (though in practice it is lower for fixed, small degrees). Since our implementation explicitly verifies any proposed bijection, false positives do not occur; verification costs are included below.

### v2

@@ Abstract
Determining whether two graphs are structurally identical is a fundamental problem with applications spanning mathematics, computer science, chemistry, and network science. Despite decades of study, graph isomorphism remains a challenging algorithmic task, particularly for highly symmetric structures. Here we introduce a new algorithmic approach grounded in spectral graph theory and discrete geometry that constructs vertex correspondences using curvature-derived, multi-scale signatures. Each vertex is assigned a canonical curvature signature built from heat-kernel coefficients organized over its breadth-first search tree, yielding a rich local invariant that can be matched globally across graphs.
Any correspondence produced by the algorithm is explicitly verified, ensuring one-sided correctness: non-isomorphic graphs are never incorrectly identified as isomorphic. Although the method does not yet guarantee success on all isomorphic inputs, we find that it correctly resolves every instance tested within deterministic polynomial time, including a broad collection of graphs known to be difficult for classical spectral and refinement-based techniques. These results demonstrate that enriched, geometry-informed spectral invariants can be substantially more discriminative than eigenvalues alone, and suggest a promising direction for practical progress on the graph isomorphism problem.
@@ 1 Introduction
In this work, we introduce a new algorithmic approach to graph isomorphism grounded in spectral geometry. Rather than exploiting combinatorial restrictions, our method draws on ideas from non-commutative geometry and spectral graph theory, using curvature derived from Laplacian spectra as a tool for symmetry breaking. The conceptual foundation of our approach is inspired by recent work on geometric features of networks based on Connes’ spectral triplet formalism [15], where distance, dimension, and curvature are recovered from spectral data via heat-kernel methods. While that framework was developed for higher-order networks and simplicial complexes, we adapt its core ideas to the setting of simple graphs and graph isomorphism.
At the core of our method is the extraction of curvature information from the short-time behavior of the heat kernel associated with the graph Laplacian. Rather than relying on eigenvalues or eigenvectors alone, we organize this curvature information into multi-scale, vertex-centered signatures that encode the local spectral geometry around each vertex. These signatures are constructed hierarchically over breadth-first search neighborhoods, enabling a local-to-global identification of vertices across graphs.
The resulting signatures are used to construct candidate correspondences between vertices of two graphs. Any correspondence produced by the algorithm is explicitly verified, ensuring one-sided correctness: the method never falsely declares two non-isomorphic graphs to be isomorphic. Consequently, any potential failure can only occur on isomorphic inputs.
We evaluate the algorithm on an extensive collection of graphs generated using standard generators, encompassing a wide range of sizes, densities, and structural properties. Remarkably, the method successfully resolves all instances tested, including graphs that are highly symmetric and traditionally challenging for classical spectral and refinement-based approaches. While this empirical success does not constitute a proof of polynomial-time solvability of graph isomorphism, it provides strong evidence that curvature-based spectral information captures structural features that are invisible to classical eigenvalue methods.
Beyond its practical performance, this work establishes a previously unexplored connection between graph isomorphism and discrete spectral geometry. It demonstrates that geometric quantities derived from spectral data, specifically curvature organized across multiple scales, can serve as effective symmetry-breaking invariants. This perspective suggests a new direction for GI algorithms, complementary to group-theoretic and combinatorial techniques, and raises the possibility that spectral geometry may play a substantive role in future progress toward polynomial-time graph isomorphism algorithms.
@@ 2.1 Spectral graph theory, Laplacians, and heat kernels
The graph Laplacian is a central object in spectral graph theory and provides a bridge between discrete graph structure and continuous analytic methods. Given an undirected graph $G=(V,E)$ with adjacency matrix $A$ and degree matrix $D$ (where $D_{ii}=\deg(i)$ and zero otherwise), the combinatorial Laplacian is defined as L=DAL=D-A. The matrix $L$ is symmetric and positive semidefinite, and its eigenvalues are real and nonnegative [14]. A fundamental property of $L$ is that it always admits the constant vector as an eigenvector with eigenvalue 0. More generally, the multiplicity of the 0 eigenvalue equals the number of connected components of $G$, with each component contributing one independent vector to the kernel of $L$ [6]. Throughout this work we focus on the combinatorial Laplacian; weighted graphs and normalized variants can be treated analogously but are not required for our analysis.
@@
These connections motivate the interpretation of $L$ as a discrete analogue of the Laplace–Beltrami operator on a Riemannian manifold. In this analogy, the heat kernel $H(t)=e^{-tL}$ plays a role analogous to its continuous counterpart. The heat kernel describes the evolution of diffusion on the graph and provides a natural multi-scale summary of its structure: small values of $t$ probe local neighborhoods, while larger $t$ reflect increasingly global connectivity. Heat-kernel-based constructions have found applications in clustering, graph comparison, and shape analysis, and they form the foundation for several modern spectral descriptors [7].
@@ 2.2 Spectral invariants and graph isomorphism
Despite this limitation, spectral methods play an important role in practical GI algorithms. Eigenvalues and eigenvectors are commonly used as preprocessing tools to partition vertices into equivalence classes, reduce symmetry, and guide subsequent combinatorial searches. However, in highly symmetric graphs, eigenvectors may become unstable or insufficiently discriminative, motivating the search for richer, vertex-level spectral invariants. More refined spectral constructions attempt to extract such invariants from Laplacian eigenfunctions rather than relying solely on eigenvalues. For example, heat-kernel-based signatures embed vertices into continuous feature spaces that capture structural information across multiple scales [17]. These methods transform isomorphism testing into a matching problem between spectral embeddings, often yielding efficient algorithms in practice even though worst-case guarantees remain elusive.
@@ 2.4 From spectra to geometry
Recent advances in spectral geometry and non-commutative geometry provide precisely such tools. By extracting geometric quantities, such as distances, dimensions, and curvature, from Laplacian spectra and heat kernels, one can associate a geometric structure to a graph that is invariant under relabeling but sensitive to subtle asymmetries. In this work, we build on these ideas to develop curvature-based, multi-scale spectral features for simple graphs and demonstrate their effectiveness for graph isomorphism testing.
@@ 3 The base algorithm
We will use Equations 2 and 5 to retrieve every node’s curvature in each graph where $u_{1}(i,i)$ will be computed from the polynomial fit of $K(i,i,t)$ versus $t$ up to $k=\text{degree(i)}$, which is the degree of node $i$. Recall that the expansion is only valid in the vicinity of $t^{+}\rightarrow 0$. Therefore, we first evaluate Equation 2 over the interval $t\in[t_{min}=1/\lambda_{max},t^{\prime}]$, where $\lambda_{max}$ is the largest eigenvalue of $L$. It defines the characteristic time scale for the heat to start diffusing over the network, while $t^{\prime}\leq 1/\lambda_{min}$, which is the time needed for the heat to diffuse over all the graphs in the network. Then a segmented linear fit of the $\log K(i,i,t)$ is applied on the data to detect a change in behavior of $K(i,i,t)$. The first point of change denotes $t_{max}$, the upper limit on the time interval over which the kernel should be followed for node $i$. Subsequently, we sample from $K(i,i,t)$, on the identified interval $t\in[t_{min}=1/\lambda_{max},t_{max}]$, a number of points equal to $\text{degree}_{i}$ and fit them against the powers of $t$ given in Equation 5 to recover the curvatures.
@@
Higher-order heat-kernel coefficients and effective curvature.
Although higher-order Minakshisundaram-Pleijel coefficients are algebraically constrained by lower-order curvature terms in the smooth Riemannian setting, they nonetheless encode additional geometric information through higher-order curvature invariants, such as contractions of the Riemann and Ricci tensors and their derivatives. In the discrete graph setting, this hierarchical dependence is considerably weaker: vertices may share identical leading-order curvature coefficients $u_{1}$ while differing at higher orders of the heat-kernel expansion. Consequently, equality of curvature does not imply equality of the full local spectral structure. In the present work, we focus on the leading-order curvature term and find it sufficient to resolve all tested instances. However, the numerical curvature values used in our algorithm are obtained by regression over finite, data-driven time windows and therefore implicitly mix contributions from multiple asymptotic orders. This implicit multiscale aggregation likely contributes to the strong empirical discriminative power observed in practice. We explicitly incorporate higher-order coefficients of the heat-kernel expansion as additional features in our proposed algorithm, which may further strengthen vertex discrimination in more challenging or adversarial instances.
Curvature versus vertex cospectrality.
It is important to emphasize that curvature, as computed in our framework, is not directly comparable to classical notions of vertex cospectrality. Vertex cospectrality is a global and exact spectral property, requiring equality of the diagonal heat kernel $K(i,i,t)$ for all times $t>0$, or equivalently equality of all moments of the local spectral measure. By contrast, our curvature estimates are explicitly scale-dependent: they are extracted from the short-time behavior of the heat kernel over a finite interval chosen to maximize geometric signal while avoiding global diffusion effects. Consequently, two vertices may share identical curvature values at short times yet diverge at longer times and therefore fail to be cospectral. Conversely, vertices that are formally cospectral may still be distinguished by finite-time geometric features arising from the interaction of multiple asymptotic orders in the numerical estimation process. For these reasons, curvature-based invariants and vertex cospectrality should be viewed as fundamentally different descriptors operating at different levels of resolution. The empirical success of our method suggests that scale-dependent geometric information can be more informative for isomorphism testing than exact but global spectral equivalence.
4 Handling degeneracies
The core algorithm described earlier relies on curvature values derived from spectral geometry to induce a vertex correspondence between two graphs. While this approach performs well in a wide range of instances, two types of degeneracies can arise in practice: graphs whose spectra are highly degenerate or insufficiently smooth, and graphs with equal curvature values across many vertices. In this section, we describe procedures that address both issues.
@@ 4.1 Stretching geometry using edge subdivisions
4.2 Handling large groups of equal curvature via vertex identification
A second source of difficulty arises when graphs possess large groups of vertices having equal curvature, e.g., large automorphism groups. In such cases, many vertices may share identical curvature values even after geometric refinement, reflecting genuine symmetries of the graph. While this does not lead to false positives, it can prevent the base algorithm from constructing a unique vertex correspondence.
To address this, we introduce an iterative vertex identification procedure that systematically breaks residual symmetries. The procedure operates as follows. At each iteration, we identify the smallest nontrivial equivalence class of vertices in the first graph whose members share the same curvature value. We perform the same identification in the second graph. If the sizes of these classes differ, the graphs are unlikely to be isomorphic and the algorithm terminates. Otherwise, we select one representative vertex from each class.
To distinguish the chosen vertices, we attach to each selected vertex a fixed number of disjoint triangles, denoted by an iteration-dependent parameter. Concretely, at iteration $t$, we attach $t$ new triangles to the selected vertex by adding $3t$ new vertices and $3t$ edges forming $t$ isolated 3-cycles incident only to that vertex. This operation creates a unique local geometric signature that is reflected immediately in curvature values, while remaining invariant under graph isomorphism when/if applied symmetrically.
The modified graphs are then reprocessed by the curvature-based algorithm, producing refined curvature values that typically split the previously degenerate equivalence classes. If necessary, the process is repeated, incrementing the number of attached triangles at each iteration. Because each step introduces strictly local modifications, the procedure is guaranteed to maintain one-sided correctness: non-isomorphic graphs are never identified as isomorphic.
@@ 4.2 Handling large groups of equal curvature via vertex identification
Together, edge subdivision and vertex identification extend the applicability of the curvature-based approach to degenerate spectral regimes and graphs with large groups of vertices having equal curvature. These mechanisms allow the algorithm to operate robustly across a broader class of inputs while preserving correctness (on tested instances) and maintaining a clear geometric interpretation.
5 The proposed complete algorithm for graph isomorphism
We now describe the full, end-to-end isomorphism test, which augments the curvature-based method with a stronger vertex invariant. For each vertex $v$ in a graph $\mathcal{G}$, we define the BFS-curvature signature of $v$, denoted $\mathrm{BCS}(v)$, to be the tuple of heat-kernel (Minakshisundaram–Pleijel) coefficients collected by breadth-first search (BFS) rooted at $v$. Concretely, $\mathrm{BCS}(v)$ consists of:
  1. 1.
    All MP coefficients $u_{k}(v,v)$ for $k=0,1,\dots,\deg(v)$ at $v$ itself;
  2. 2.
    Then, for each distance-1 neighbor $w$ of $v$, its MP coefficients $u_{k}(w,w)$ (sorted in nondecreasing lexicographic order of their tuples);
  3. 3.
    Then similarly the sorted lists of MP coefficients of all vertices at distance 2 from $v$;
  4. 4.
    And so on, level by level, following the BFS tree rooted at $v$ until all vertices of the graph are included.
Two vertices (possibly in different graphs) are called BFS-curvature equivalent if their BFS-curvature signatures are identical. This signature encodes the full local spectral geometry around a vertex, not just its leading-order curvature. As noted previously, vertices can share the same curvature $u_{1}$ yet differ at higher orders of the heat-kernel expansion; the BFS-curvature signature remedies this by including all such coefficients in a canonical order.
The complete isomorphism algorithm proceeds as follows:
  1. 1.
    Signature computation. For each original vertex in both input graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, compute its BFS-curvature signature $\mathrm{BCS}(v)$ by performing a BFS from $v$ and collecting the MP coefficients at each layer as described above.
  2. 2.
    Equivalence classes. Partition the vertices of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ into equivalence classes according to their signatures. If at any stage an equivalence class in $\mathcal{G}_{1}$ has a different size than the corresponding class in $\mathcal{G}_{2}$, the graphs are declared not to be isomorphic and the algorithm terminates. Otherwise, each class contains vertices that are pairwise BFS-curvature equivalent.
  3. 3.
    Refinement (if needed). If every equivalence class consists of a single vertex, the vertices are uniquely distinguished and we proceed to matching. Otherwise, apply the geometric refinement steps as described in the previous section: subdivide edges and attach distinguishing gadgets to break remaining symmetries. In practice we apply the same edge-subdivision and vertex-augmentation operations to both graphs in parallel. Edge subdivision “stretches” the geometry and spreads the spectrum, while attaching local triangles creates unique curvature signatures at selected vertices. After each refinement, recompute all BFS-curvature signatures and re-partition into classes. Repeat this process until all original vertices are distinguished.
  4. 4.
    Signature matching. Once every vertex in each graph has a unique BFS-curvature signature, we form the final correspondence by sorting and pairing. Concretely, sort the vertices of $\mathcal{G}_{1}$ in lexicographic order of $\mathrm{BCS}(v)$ and do the same for $\mathcal{G}_{2}$. This yields a one-to-one mapping between vertices of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. In other words, corresponding entries in the sorted signature lists define the vertex bijection.
Local versus global isometry.
The BFS-curvature signature can be viewed as a discrete analogue of the local curvature profile of a point on a Riemannian manifold. In differential geometry, the full collection of curvature invariants (and their derivatives) at a point determines the local metric up to isometry (Cartan–Ambrose–Hicks theorem). Similarly, $\mathrm{BCS}(v)$ encodes the geometry of the metric-ball around $v$ in the graph. In fact, this idea parallels the heat-kernel signature (HKS) used in shape analysis [18], which restricts the continuous heat kernel to the diagonal and captures the local diffusion geometry at a point. Matching BFS-curvature signatures across two graphs is therefore analogous to constructing a global isometry: if every vertex in $\mathcal{G}_{1}$ has an identical signature to some vertex in $\mathcal{G}_{2}$, then the two graphs share the same discrete metric structure. In effect, sorting and aligning the signatures from $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ “enforces” a discrete global isometry between the graphs.
In summary, the complete algorithm uses BFS-curvature signatures to canonically label every vertex with a rich spectral invariant. Vertices are matched by simply comparing these signatures. This global matching of local signatures yields a deterministic correspondence: the graphs are declared isomorphic if and only if the sorted lists of BFS-curvature signatures coincide. The final output is thus a vertex bijection defined by this signature matching. Notably, other approaches to discrete curvature (e.g. Ollivier–Ricci or Forman curvature) also link local geometry to global structure, but our method is based on the heat-kernel expansion and BFS layering instead [16, 10].
The idea of using layered spectral signatures builds on prior work in spectral graph invariants and heat-kernel methods (e.g. [17, 7, 18, 16, 10]). In particular, [17] demonstrated that heat kernel signatures can distinguish vertices in isomorphic graphs, and we extend this by organizing higher-order heat-kernel coefficients in a BFS hierarchy. This yields a complete matching algorithm analogous to a discrete version of global isometry in geometry.
@@ 6 Computational complexity
We analyze the running time of three progressively stronger components of our approach: (i) the base curvature–matching algorithm, (ii) the degeneracy-handling mechanisms, and (iii) the full BFS-curvature–signature algorithm. Throughout, let $n=|V|$ and $m=|E|$ denote the number of vertices and edges of the input graphs. We assume standard RAM-model arithmetic on $O(\log n)$-bit numbers and treat polynomial regression on $O(d)$ points as $O(d^{3})$ in the worst case (though in practice it is significantly lower for small, fixed degrees). Since our implementation explicitly verifies any proposed vertex correspondence, false positives do not occur; verification costs are included in the bounds below.